OpenStudy (liv33):

Solve for the variable: 20 - 3h = 2

Mehek (mehek14):

Subtract by 20 -3h = 2 - 20 2 - 20 = ?

OpenStudy (liv33):

-18

Mehek (mehek14):

Divide by -3 -18/-3

OpenStudy (liv33):

6

Mehek (mehek14):

h = 6

OpenStudy (liv33):

Thanks :)

Mehek (mehek14):

You're welcome :D